OpenStudy (anonymous):

WILL GIVE MEDAL Find the vertex of this parabola. y=-2x^2-8x+7.

OpenStudy (anonymous):

y=2x^2-8x+7 Complete the square and format as a parabola x^2 -4 x = 1/2 y -7/2 Complete the square of the x terms by adding ( -4/2 )^2 = 4 to both sides + x^2 -4 x +4 = 1/2 y -7+4 (x-2)^2 = 4 (1/8) (y-6) The focal parameter, p = 1/8 Since p > 0 then the parabola opens up. The vertex of the parabola is (h,k): (h,k) = (+2, +6)

OpenStudy (anonymous):

no its -2,15

OpenStudy (anonymous):

ok